OpenStudy (anonymous):

In a certain medium, sound travels at a speed of 7189 feet per second. How many meters per second does sound travel in this medium? Round to the nearest tenth.

OpenStudy (anonymous):

\[\frac{ 7189 feet }{ 1 \sec } \times \frac{ 12 inches }{ 1 foot } \times \frac{ 2.54 centimeters }{ 1 inch } \times \frac{ 1 meter }{ 100 centimeters }\] you'll want to just multiply through these conversion factors. notice how all the units cancel out except for meter/sec and the end

OpenStudy (anonymous):

at the end*

OpenStudy (anonymous):

ok thanks